OpenStudy (anonymous):

Please Solve lim x-0 cotx/lnx using L,hopital's rule.... anybody help me ?

hartnn (hartnn):

so did you take the derivative of numerator and denominator separately ?? what did u get ?

hartnn (hartnn):

derivative of cot x is ?? derivative of ln x is ?

OpenStudy (dls):

In L-Hospital Rule we generally take the derivative of numerator and denominator separately. You might need some references. \[\Large \frac{d}{dx}(\cot x)=?\] \[\Large \frac{d}{dx}(\ln x)=?\]

OpenStudy (anonymous):

-csc^2x and 1/x

hartnn (hartnn):

so you have -x csc^2 x right ?? how do u convert csc x to sin x ?>

OpenStudy (dls):

nt rly reqd :o

hartnn (hartnn):

you will need this : lim x->0 sin x/x =1

hartnn (hartnn):

cosec 0 is infinity, right ? 0*infinity is still not defined

OpenStudy (dls):

\[\Large \frac{-cosec^2 x}{\frac{1}{x}} = \lim_{x \rightarrow 0}- x cosec^2 x\] anything into 0 is 0

OpenStudy (dls):

its 0 only...

hartnn (hartnn):

anything except infinity **

hartnn (hartnn):

limiting value is 0 only but there's a mathametical way to prove that, so why not

OpenStudy (dls):

im out here

OpenStudy (anonymous):

thanks DLS and hartnn

hartnn (hartnn):

you got how to solve ?

hartnn (hartnn):

and sure, you're welcome ^_^

OpenStudy (anonymous):

yes i'll solve it now..

OpenStudy (anonymous):

isn't this function divergent?